# Scaling behavior of dissipative systems with imaginary gap closing

### v1

@@ III Scaling behavior of imaginary gap-closing systems with a trivial point-gap topology
where $t_{0}$ and $t_{1}$ are the intracell and intercell hopping amplitudes, respectively, $\gamma$ is the onsite dissipation on sublattice B, and H.c. denotes the Hermitian conjugate. By applying a Fourier transformation and making the substitution $\beta\rightarrow e^{ik}$, we obtain the generalized non-Hermitian Bloch Hamiltonian:
@@
where $\alpha^{\prime}=-\frac{g_{+}^{AA}\left(-1\right)\Gamma\left(1/4\right)}{8\pi i\left(-E_{+}^{\left(4\right)}\left(-1\right)/24\right)^{1/4}}$. As a result, the asymptotic scaling of the local Green’s function $G_{A}\left(x_{0};t\right)$ at these critical values $\left|t_{0}\right|=\left|2t_{1}\right|$ is
@@ IV Scaling behavior of imaginary gap-closing systems with a nontrivial point-gap topology
For nonzero $t_{0}$, the solutions of this equation do not lie in the unit circle $\left|\beta\right|=1$, implying that the corresponding energies are not part of the PBC spectrum. In contrast, the endpoints of the OBC spectra are attained for values of $\beta$ on the GBZ that correspond to the saddle points of $E_{1,\pm}\left(\beta\right)$, as shown in Figs. 3(a) and 3(b). For a general model, not all saddle points are located on the GBZ [86, 87]. Instead, they generally reside on the auxiliary generalized Brillouin zone (aGBZ) [88, 89].
@@
$G_{A}\left(x_{0};t\right)\approx$ $\sum_{n=\pm}\frac{g_{1,n}}{2\pi i\beta_{s}}g_{1,n}^{AA}\left(\beta_{s}\right)e^{-iE_{1,n}\left(\beta_{s}\right)t}$ (22)
$\sim$ $e^{\mathrm{Im}\left(S\right)t},$
@@ Acknowledgements.
The authors thank Xinyuan Gao for helpful discussion. This work is supported by the Quantum Science and Technology-National Science and Technology Major Project (2024ZD0300600). We acknowledge financial support from the National Natural Science Foundation of China under Grant No. 92565105 and 12204395, Hong Kong RGC No. 14301425, No. 24308323, and No. C4050-23GF, the Space Application System of China Manned Space Program, and CUHK Direct Grant. Jinghui Pi also acknowledges support from the Postdoctoral Fellowship Program of CPSF under Grant No. 2025M773387.

### v2

@@ III Scaling behavior of imaginary gap-closing systems with a trivial point-gap topology
where $t_{0}$ and $t_{1}$ are the intracell and intercell hopping amplitudes, respectively, $\gamma$ is the onsite dissipation on sublattice B, and H.c. denotes the Hermitian conjugate. By applying a Fourier transformation and making the substitution $e^{ik}\rightarrow\beta$, we obtain the generalized non-Hermitian Bloch Hamiltonian:
@@
where $\alpha^{\prime}=-\frac{g_{+}^{AA}\left(-1\right)\Gamma\left(1/4\right)}{8\pi i\left[-E_{+}^{\left(4\right)}\left(-1\right)/24\right]^{1/4}}$. As a result, the asymptotic scaling of the local Green’s function $G_{A}\left(x_{0};t\right)$ at these critical values $\left|t_{0}\right|=\left|2t_{1}\right|$ is
@@
The above discussion does not always hold for imaginary gap-closing systems with a trivial point-gap topology. In fact, at these imaginary gap-closing points $\beta_{0}$, we have $\operatorname{Im}\partial_{\beta}E(\beta)|_{\beta=\beta_{0}}=0$. Hence, if $\operatorname{Re}\partial_{\beta}E(\beta)|_{\beta=\beta_{0}}=0$, these points are also saddle points and correspond to the endpoints of the energy spectrum $E(\beta)$. However, it is also possible that $\operatorname{Re}\partial_{\beta}E(\beta)|_{\beta=\beta_{0}}\neq 0$, which means that $\beta_{0}$ is not a saddle point of $E(\beta)$. In this case, the energy spectrum $E(\beta)$ is tangent to the $\operatorname{Re}E$ axis at $\beta_{0}$ in the complex plane.
@@ IV Scaling behavior of imaginary gap-closing systems with a nontrivial point-gap topology
For nonzero $t_{0}$, the solutions of this equation do not lie in the unit circle $\left|\beta\right|=1$, implying that the corresponding energies are not part of the PBC spectrum. In contrast, the endpoints of the OBC spectra are attained for values of $\beta$ on the GBZ that correspond to the saddle points of $E_{1,\pm}\left(\beta\right)$, as shown in Figs. 3(a) and 3(b). For a general model, not all saddle points are located on the GBZ [86, 87]. In fact, the corresponding characteristic equation $\det\left[h\left(\beta\right)-E\right]=0$ gives an exact relation between $E$ and $\beta$. The saddle points are independent of the boundary conditions and are uniquely determined by the equation $\partial_{\beta}E(\beta)=0$, indicating that they correspond to double roots of the characteristic equation. Hence, these saddle points typically reside on the auxiliary generalized Brillouin zone (aGBZ) [88, 89, 90].
@@
$G_{A}\left(x_{0};t\right)\approx\sum_{n=\pm}\frac{g_{1,n}^{AA}\left(\beta_{s}\right)}{2\pi i\beta_{s}}e^{-iE_{1,n}\left(\beta_{s}\right)t}\sim e^{\mathrm{Im}\left(S\right)t},$ (22)
@@
At the end of this section, we briefly examine the imaginary gap closing of the OBC spectrum. In a general non-Hermitian system, a point-gap nontrivial PBC $E_{PBC}$ spectrum encloses the corresponding OBC spectrum $E_{OBC}$, resulting in the relation $\max\left[\operatorname{Im}E_{OBC}\right]\leq\max\left[\operatorname{Im}E_{PBC}\right]$, where equality generally cannot be attained [39]. Consequently, for dissipative systems featuring the NHSE, the OBC spectrum typically does not satisfy the imaginary gap-closing condition. To enable the OBC spectrum of the Hamiltonian (17) to satisfy $\max\left[\operatorname{Im}E_{OBC}\right]=0$, additional gain can be introduced to the entire system. This shifts the endpoints of the OBC spectrum onto the $\operatorname{Re}E$ axis. In this case, the saddle-point method remains applicable to the OBC system. However, the scaling behavior of the local Green’s function becomes dependent on the site $x_{0}$: $G_{A}\left(x_{0};t\right)$ exhibits a $t^{-1/2}$ scaling in the bulk but a $t^{-3/2}$ scaling at the edge [87].
@@ Acknowledgements.
The authors thank Xinyuan Gao for helpful discussion. This work is supported by the Quantum Science and Technology-National Science and Technology Major Project (2024ZD0300600). We acknowledge financial support from the National Natural Science Foundation of China under Grant No. 92565105 and 12204395, Hong Kong RGC No. 14301425, No. 24308323, and No. C4050-23GF, Guangdong Provincial Quantum Science Strategic Initiative GDZX2404004, GDZX2505005, the Space Application System of China Manned Space Program, and CUHK Direct Grant. Jinghui Pi also acknowledges support from the Postdoctoral Fellowship Program of CPSF under Grant No. 2025M773387.